OpenStudy (anonymous):

Find the inverse of the following matrix, if it is defined. 3 2 3 -2

OpenStudy (anonymous):

Find the determinant first. \[\det \left|\begin{matrix}a & b \\ c & d\end{matrix}\right|=ad-bc\]